# Shear stresses in fluid and solid membranes with bending elasticity

### v1

@@ II Fluidity and its malcontents
$-\nabla_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(SNa)}}\cdot{\mathbf{n}}=\tfrac{1}{2}\nabla_{\alpha}\nabla\gamma\left(A^{\beta\gamma}b_{\beta}^{\alpha}\right)-2H\sigma_{\text{\tiny(SNa)}}\,.$ (14)
VI Discussion
@@
The existence or absence of shear stresses is a consequence of a constitutive choice, and the elegance and ease of definition of a geometric energy belies its complex response. Quadratic-curvature energies, such as Helfrich, are such that curvatures give rise to tangential stresses, and these are in general anisotropic. Energies quadratic in the product of stretch and curvature, such as Atluri/SN, simply do not have this coupling property. This result is only indirectly related to the requirement of fluidity, which restricts the quantities that can be used to the Jacobian determinant and the geometric curvature invariants [7]. It may be possible to construct a fluid-like energy quadratic in the product of stretch and curvature that consists only of a term analogous to the curvature determinant; a term analogous to the square of the curvature trace seems to be precluded, although this remains an open question.
@@
While there is a stark qualitative distinction between these models, the differences are nonetheless quantitatively only of the order of the in-plane strain. This is small for stiff solids, for which there is a penalty for any type of strain, and with metric constraints the physical distinctions disappear altogether. However, constraining only the area does not limit the size of shear strains. In such cases, the differences in shear stress are on the order of the curvature, comparable to other stresses generated by bending.
The present results regarding stresses and bend-stretch coupling are likely most relevant to the modeling of elastic capsules in viscous flows [52], which sometimes combines an area-constrained Helfrich energy with an additional penalty on surface stretching. Within that literature, a variety of other models have been employed, including the complete neglect of bending resistance. Pozrikidis [53] posits a linear moment-curvature relation without coupling to in-plane stresses, assuming “sufficiently small bending deformations”. Without this restriction, the prescription is not consistent with variational derivation from a quadratic-curvature (or perhaps any) energy. He contrasts this with Zarda and co-workers’ [19] use of a Reissner-like axisymmetric shell bending measure to model red blood cells, and further recognizes that this measure is related to dilation invariance. We note that a bending energy quadratic in the Reissner quantities would be akin to an Atluri/SN energy, and lead to a linear, decoupled constitutive response [12, 13, 14, 36, 37].
@@
Our concern has been with elucidating the existence and origin of certain terms, and the structure of the equations. We remain agnostic on the value of any of these models in accurately representing any type of physical membrane, or their possible derivation from microscopic physics.
VII Conclusions
Shear stresses in membranes with bending elasticity were discussed in the context of a few simple and common fluid and solid models, alongside the structure of the momentum and pseudomomentum balances, and some commentary on the literature.
@@ Acknowledgments
We thank TJ Healey, DJ Steigmann, and particularly E Vitral, for discussions. JAH was supported by U.S. National Science Foundation grant CMMI-2001262.

### v2

@@ II Fluidity and its malcontents
$-\nabla_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(SNa)}}\cdot{\mathbf{n}}=\tfrac{1}{2}\nabla_{\alpha}\nabla_{\gamma}\left(A^{\beta\gamma}b_{\beta}^{\alpha}\right)-2H\sigma_{\text{\tiny(SNa)}}\,.$ (14)
The rudimentary example in the following Section VI highlights a difference between (14) and the normal projection of (12), also making use of results from the previous Section IV.
VI Example: Unidirectional bending
A rudimentary example, using results from Sections IV and V, serves to illustrate important differences between the fluid Helfrich (1) and solid area-constrained Atluri/SN bending (3) models. For simplicity of calculation, we consider a patch of membrane uniform in curvature and metric, such that terms involving derivatives of these vanish. The condition on curvature is actually rather restrictive, leading to discrete special cases of unbent planar, isotropically bent spherical, or unidirectionally bent cylindrical patches, with only the last being anisotropic in curvature. The cylindrical example includes the physical case of the tether mentioned in Section II, and displays the relative complexity of the Helfrich response and simplicity of the Atluri/SN response.
The multipliers contributing to the isotropic part of the tractions can be specified by balancing the normal component of the divergence of stress from (12) or (14) against a suitably nondimensionalized pressure difference $\Delta P$ between the two sides of the membrane. For a uniform patch, we obtain $\sigma_{\text{\tiny(H)}}=H^{2}-K+\Delta P/2H$ and $\sigma_{\text{\tiny(SNa)}}=\Delta P/2H$.
The tractions $\nu_{\alpha}{\mathbf{f}}^{\alpha}$ at each point on the boundary of a deformed patch can be considered in terms of their components on the local orthogonal unit triad consisting of the surface normal ${\mathbf{n}}$ and two tangential directions, the tangent normal ${\bm{\nu}}$ $(=\nu_{\alpha}{\mathbf{a}}^{\alpha})$ and tangent tangent ${\bm{\tau}}$. Consider a uniform cylindrical patch, curved in a single direction. From the stresses (7) and (9), we see that the normal component $\nu_{\alpha}{\mathbf{f}}^{\alpha}\cdot{\mathbf{n}}=0$ for either model, and we further obtain $\nu_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(H)}}\cdot{\bm{\nu}}=-Hb_{\nu\nu}+\sigma_{\text{\tiny(H)}}+H^{2}$, $\nu_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(H)}}\cdot{\bm{\tau}}=-Hb_{\nu\tau}$, $\nu_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(SNa)}}\cdot{\bm{\nu}}=\sigma_{\text{\tiny(SNa)}}$, and $\nu_{\alpha}{\mathbf{f}}^{\alpha}_{\text{\tiny(SNa)}}\cdot{\bm{\tau}}=0$ (with these local orthonormal $\nu$-$\tau$ coordinates, the index type is immaterial, and $2H=b_{\nu\nu}+b_{\tau\tau}$). This situation is shown in Figure 1; not shown are pressures normal to the membrane on either side, and moments around the tangent tangent, also necessary to sustain equilibrium.
Figure 1: Components of boundary tractions $\nu_{\alpha}{\mathbf{f}}^{\alpha}$ on a uniform patch of unidirectionally-curved membrane corresponding to two area-constrained bending energy models, (a) Helfrich and (b) Atluri/SN, shown at one location on the boundary and expressed in local orthonormal coordinates there. Uniformity eliminates terms involving derivatives, and also components normal to the membrane, leaving only the in-plane tangent normal $\nu$ and tangent tangent $\tau$ (shear) components, with simplified forms. Not shown are pressures normal to the patch on either side, and moments around the tangent tangent, also necessary to sustain equilibrium. The fluid Helfrich tractions are anisotropic, while the solid Atluri/SN is tangentially isotropic.
The Atluri/SN solution is particularly simple. There is a tangentially isotropic (pressure-like) stress. Furthermore, if there is no difference $\Delta P$ in external normal pressures, then $\sigma_{\text{\tiny(SNa)}}$ must be zero (as $H$ is nonzero), and there are no tangential tractions at all. The relation between moments and bending is decoupled from the in-plane stresses. The Helfrich solution is more complex. The tractions feature anisotropic contributions, whose appearance in the tangent normal or tangent tangent (shear) directions depends on the local orientation of the boundary. The curvatures $b_{\nu\nu}$ and $b_{\nu\tau}$ vary around the boundary. At four points corresponding to two principal directions, $b_{\nu\nu}$ takes on its extreme values and $b_{\nu\tau}$ vanishes. A similar statement can be made about the in-plane stresses themselves. For such details of the behavior of symmetric two-dimensional tensors, we refer the interested physicist to the exposition of Mohr’s circle in an undergraduate engineering text such as [52]; however, to reproduce an analogously simple picture of a stressed small square of material for an arbitrary curved surface, the derivative terms in the tractions, including normal components, need to be considered.
Returning to the case of the pulled membrane tether, we note that Powers and co-workers [23], like others since, assumed that such experimentally realized tethers are cylindrical, with no pressure difference between the two sides of the membrane. They pointed out that the anisotropy of the Helfrich stress is necessary to construct this solution while satisfying a boundary condition corresponding to a pulling force. For a cylindrical shape ($K=0$) and no pressure difference ($\Delta P=0$), we find $\sigma_{\text{\tiny(H)}}=H^{2}$ in the Helfrich solution. Inserting values, we observe that there is no traction on portions of the boundary whose tangent normals are in the bending direction, but on portions of the boundary in the orthogonal direction, a tangent normal component of $2H^{2}$ is required to sustain equilibrium. There is an applied force in the unbent (axial) direction, while the circumferential (hoop) stress is zero. On the other hand, the Atluri/SN bending energy does not admit such solutions, requiring either a pressure difference, or a non-cylindrical shape. For this bending energy, we find $\sigma_{\text{\tiny(SNa)}}=0$, which precludes the existence of any applied force; the cylindrical tether with an axial force and no pressure difference is not a solution. More generally, the applied force and pressure difference are directly linked.
VII Discussion and Conclusions
@@
The existence or absence of shear stresses is a consequence of a constitutive choice, and the elegance and ease of definition of a geometric energy belies its complex response. Quadratic-curvature energies, such as Helfrich, are such that curvatures give rise to tangential stresses, and these are in general anisotropic. Energies quadratic in the product of stretch and curvature, such as Atluri/SN, simply do not have this coupling property. This result is only indirectly related to the requirement of fluidity, which restricts the quantities admissible in an energy to the Jacobian determinant and the geometric curvature invariants [7]. It may be possible to construct a fluid-like energy quadratic in the product of stretch and curvature that consists only of a term analogous to the curvature determinant; a term analogous to the square of the curvature trace seems to be precluded, although this remains an open question.
@@
While there is a stark qualitative distinction between these models, the differences are nonetheless quantitatively only of the order of the in-plane strain. This is small for stiff solids, for which there is a penalty for any type of strain, and with metric constraints the physical distinctions disappear altogether. However, constraining only the area does not limit the size of shear strains. In such cases, the differences in shear stress are of the order of the curvature, comparable to other stresses generated by bending.
@@
The present results regarding stresses and bend-stretch coupling are likely relevant to the modeling of elastic capsules in viscous flows [53], which sometimes combines an area-constrained Helfrich energy with an additional penalty on surface stretching. Within that literature, a variety of other models have been employed, including the complete neglect of bending resistance. Pozrikidis [54] posits a linear moment-curvature relation without coupling to in-plane stresses, assuming “sufficiently small bending deformations”. Without this restriction, the prescription is not consistent with variational derivation from a quadratic-curvature (or perhaps any) energy. He contrasts this with Zarda and co-workers’ [19] use of a Reissner-like axisymmetric shell bending measure to model red blood cells, and further recognizes that this measure is related to dilation invariance. We note that a bending energy quadratic in the Reissner quantities would be akin to an Atluri/SN energy, and lead to a linear, decoupled constitutive response [12, 13, 14, 36, 37].
Our concern has been with elucidating the existence and origin of certain terms, and the structure of the equations. Whether these models accurately represent any type of physical membrane, or are derivable from microscopic physics, are separate issues. If the common assumptions surrounding tether formation are correct, so that shear stresses are required, another prevailing assumption is that these do not come from in-plane shear elasticity. However, as summarized in [55], Calladine and co-workers have suggested adding a small in-plane shear modulus, and coupling terms present in a “thick” shell with rest curvature, to resolve discrepancies between predictions derived using the Helfrich energy, with its bend-stretch coupling, and those derived using a direct model with linear, decoupled response, like that of Pozrikidis above. The decoupled direct model fails to explain cylindrical tether pulling and oblate buckling phenomena. The restrictions on the form of a fluid membrane energy are based on the presumption of in-plane fluidity, not three-dimensional fluidity of a thin bulk film of fluid [30]. A simple film of fluid would not have a bending energy, including aspects of a static bending response such as transverse shear stresses. The Helfrich model assumes that the membrane’s material properties endow it with an elastic bending energy while still retaining in-plane fluidity. The biophysics literature provides some suggestions of microscopic origins for bending resistance of a bilayer [56]. Red blood cells (a type of capsule) have a small shear modulus, associated with an additional spectrin-actin network [57, 56]. However, Waugh [58] found that red blood cells with spectrin abnormalities also have a reduced bending stiffness correlating with their reduced shear stiffness, implying a significant solid contribution to their bending energy.
The Helfrich energy has been successfully applied to explain static and dynamic behaviors of structures formed from closed lipid bilayer membranes. The only problems with its use that we warn against are the misinterpretation of shear stresses, and the “validation” of analytical results from Helfrich using SN-type numerics whose continuum limit is a different energy with different qualitative behavior.
@@ Acknowledgments
We thank TJ Healey, DJ Steigmann, and particularly E Vitral, for discussions prior to the initial draft, and M. Deserno and T. Powers for discussions afterwards. JAH was supported by U.S. National Science Foundation grant CMMI-2001262.